HIGH-RESOLUTION PFI-ZEKE PHOTOELECTRON SPECTROSCOPY OF Cl_{2}: THE GROUND (X^{+} ^{2}Π_{g}) AND FIRST EXCITED (A^{+} ^{2}Π_{u}) ELECTRONIC STATES OF Cl_{2}^{+}

\underline{SANDRO MOLLET}, and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang Pauli-Strasse 10, 8093 Zürich, Switzerland.

Recently, two studies have been devoted to the low-lying electronic states of Cl_{2}^{+}, one by PFI-ZEKE photoelectron (PE) spectroscopy of the X^{+} ← X transition^{a} and one by laser-induced-fluorescence spectroscopy of the A^{+} → X^{+} band system^{b}. To complement the information available on the X^{+} and A^{+} electronic states of Cl_{2}^{+}, we have recorded partially rotationally resolved single-photon PFI-ZEKE PE spectra of the X^{+} ^{2}Π_{g,i} ← X^{1}Σ_{g}^{+} and A^{+} ^{2}Π_{u,i} ← X^{1}Σ_{g}^{+} (i = 3/2, 1/2) photoionizing transitions of Cl_{2} in the wavenumber ranges 92 500−96 500 cm^{−1} and 106 750−115 500 cm^{−1}. These regions correspond to transitions to low-lying vibrational levels of the X^{+} state with v^{+} = 0 − 5 and to transitions to vibrational levels of the X^{+} state with v^{+} = 25 − 45 and vibrational levels of the A^{+} state with v^{+} ≤ 7. The analysis of the rotational structure and the isotopic shifts of these spectra has enabled the derivation of an improved value of the first adiabatic ionization energy (92 647.7±0.3 cm^{−1} for ^{35}Cl_{2}). In combination with measurements of ion-pair states with n ≈ 1800^{c}, new values for the dissociation energies D_{0} of Cl_{2} and Cl_{2}^{+} could also be derived. The potential energy function of the X^{+} state of Cl_{2}^{+} was determined in a least-squares fitting procedure. Spin-orbit splittings were derived for many vibrational levels of the X^{+} and A^{+} states. Combining our results with other results^{b,}^{d}^{,}^{e}^{,}^{f}, several low-lying vibrational levels of the upper spin-orbit component of the A^{+} state could be assigned for the first time. The PFI-ZEKE PE spectra also contain a series of as yet unassigned lines and reveal numerous perturbations.

aJ. Li, Y. Hao, J. Yang, C. Zhou, and Y. Mo, J. Chem. Phys. 127, 104307 (2007).

bM. A. Gharaibeh, and D. J. Clouthier, 66^{th} Int. Symposium on Molecular Spectroscopy, Book of Abstracts, p. 169 (2011).

cS. Mollet, and F. Merkt, Phys. Rev. A 82, 032510 (2010).

dR. P. Tuckett, and S. D. Peyerimhoff, Chem. Phys. 83, 203 (1984).

eL. Wu, X. Yang, Y. Guo, L. Zheng, Y. Liu, and Y. Chen, J. Mol. Spect. 230, 72 (2005).

fH. van Lonkhuyzen, and C. A. de Lange, Chem. Phys. 89, 313 (1984).